OpenStudy (anonymous):

show that the equation 11x + 8y + 17 = xy has 16 solutions and also list the solutions. pls help me

OpenStudy (anonymous):

pls answer any1

OpenStudy (anonymous):

checking right now, will post when answered

OpenStudy (anonymous):

ok thanks but not forget to answer me bye

OpenStudy (anonymous):

do my best

OpenStudy (anonymous):

they do not have to be integer values I pressume

OpenStudy (anonymous):

there are more than 16 solutions. 11x+17=xy-8y 11x+17=y(x-8) y=(11x+17)/(x-8) plug in any x except for x=8 and you will get a corresponding y value. All of the solutions look like this.

OpenStudy (anonymous):

example: x=1 (11+17)/(1-8)=-4 11(1)+8(-4)+17=11-32+17=-4=(-4)(1)=yx

OpenStudy (anonymous):

so (1,-4) is a solution. you can find the 16 you need this way

OpenStudy (anonymous):

rsvitale is correct but you could be there all day plugging in values of x You might notice that the odd values of x between 1 and including 7 work giving (x,y) = (1, -4) (3, -10) (5, -24) and (7, -94) You have 8 solutions now. A way forward from here would possibly be to select useful negative x values.

OpenStudy (anonymous):

...good to try these ideas but the following solutions are what you need (x,y) = (9, 116) (13, 32) (15, 26) and (25, 12)